Hamed Holding is the sole producer and seller of widgets in the economy. Widgets can be produced with a constant returns technology at a cost of 20 OR. The demand for widgets is such that the price elasticity of demand is 1.5.
20 a. What price would Aziz Holding charge if it were to maximize its profits?
10 b. What would be the market price if this were a competitive market (with many other firms all having the same technology)?
1
Expert's answer
2014-05-14T11:11:53-0400
If ATC = MC = 20 and Ed = 1.5, the monopolist’s pricing rule as a function of the elasticity of demand for its product is:
(P - MC)/P = 1/Ed
1 - 20/P = 2/3
20/P = 1/3
P = $60.

For competitive market MR = MC = P = $20, so the price will be 3 times lower and no one will have profit.